# Attribute Structuring Improves LLM-Based Evaluation of Clinical Text Summaries

### v1

@@ Abstract
Summarizing clinical text is crucial in health decision-support and clinical research. Large language models (LLMs) have shown the potential to generate accurate clinical text summaries, but still struggle with issues regarding grounding and evaluation, especially in safety-critical domains such as health. Holistically evaluating text summaries is challenging because they may contain unsubstantiated information. Here, we explore a general mitigation framework using Attribute Structuring (AS), which structures the summary evaluation process. It decomposes the evaluation process into a grounded procedure that uses an LLM for relatively simple structuring and scoring tasks, rather than the full task of holistic summary evaluation. Experiments show that AS consistently improves the correspondence between human annotations and automated metrics in clinical text summarization. Additionally, AS yields interpretations in the form of a short text span corresponding to each output, which enables efficient human auditing, paving the way towards trustworthy evaluation of clinical information in resource-constrained scenarios. We release our code, prompts, and an open-source benchmark at https://github.com/microsoft/attribute-structuring/.
[…]
(Hide=100, Hang=true, Progressive=2ex, Style*$\bullet$ , Style2*$\circ$)
Attribute Structuring Improves LLM-Based Evaluation
of Clinical Text Summaries
Zelalem Gero, Chandan Singh, Yiqing Xie, Sheng Zhang, Tristan Naumann, Jianfeng Gao, Hoifung Poon Microsoft Research
@@ 1 Introduction
Automatic medical text summarization has emerged as a critical task for real-world impact in natural-language processing, amidst widespread physician fatigue and burnout. Meanwhile, pre-trained large language models (LLMs) have shown impressive proficiency in a range of complex natural language processing tasks [1, 2, 3], including in text summarization [4, 5]. However, evaluating these summarizations is challenging, as they rely on human effort to track different attributes and facts.
Here, we propose using LLMs for evaluating medical text summaries, guided by a simple structuring process. LLMs have previously been used to provide practical evaluations in a variety of evaluation settings [6, 7]. However, these evaluations have been imperfect: yielding relatively low correlation with human judgments, showing bias [8, 9], or more generally struggling with long-form nuanced evaluation, as is often the case in summarization [10]. Structuring the scoring process can help to ground the evaluation, and help an LLM decompose a complex evaluation process into simpler, more accurate steps.
We introduce a simple structuring process, which we call Attribute Structuring (AS) (Fig. 1). Given a groundtruth summary and a summary to be evaluated, we leverage an LLM to score a summary through prompting. Rather than directly asking the LLM for a score, we ask it to provide a score for each pair of attributes, one from the groundtruth and another from the summary to be evaluated. We extract 17 attributes derived from a clinical ontology for each summary; the attributes are shown to define the elements of quality in the context of a clinical discharge summary in a large scale study [11]. Unlike previous methods such as FActScore [12], this does not require a reliable knowledge source to verify each attribute, but instead relies on the pre-specified ontology to capture relevant aspects of a summary. DocLens  [13], similar to our work, performs a fine-grained evaluation of medical text using various generation models and show that such scoring has higher agreement with the judgements of medical experts. However, DocLens relies on the model to generate the important attributes rather than a clinical ontology.
@@
We perform experiments on clinical summarization tasks using discharge notes from MIMIC-III [14] and various LLMs, including GPT-4 [3], GPT-3.5 [15], and LLaMA-2 [2]. We find that that Attribute Structuring helps reduce the gap between automated metrics and human annotators. In addition, AS enables auditing the evaluation for a particular summary, by providing grounding for each attribute, in the form of a short text span in the input.
@@ 2.1 Methods: Attribute Structuring
Fig. 1 shows the various steps of the Attribute Structuring (AS) pipeline. The pipeline takes in two summaries: a groundtruth summary and a generated summary (which we are trying to evaluate). Additionally, it uses a set of attributes that are important for the task at hand. These attributes comes from an ontology that utilizes domain knowledge to define the elements of quality, in our case the context of a clinical discharge summary in a large scale study [11], e.g. Admission diagnosis or Goals of care.
Each step in the pipeline is achieved by prompting an LLM. In the first step, the Attribute Structuring pipeline first extracts text for each attribute in the ontology from both summaries. Then, for each pair of extracted attributes, it generates a similarity score by prompting an LLM to compare them on a scale of 1–5 (all prompts are provided in Appendix A). Finally, an average of these scores is returned as the evaluation score for the generated summary (usually after normalizing to range from 0–100) .
@@ 2.2 Experimental setup
Dataset and benchmark
We construct a benchmark consisting of discharge summaries for 300 randomly selected MIMIC-III notes  [14] and release scripts to process and use this data. To be fair to LLMs with different context lengths, we select MIMIC notes with no more than 4,000 tokens (measured with tiktoken’s p50_base tokenizer [^1] https://github.com/openai/tiktoken). As it is derived from MIMIC, this benchmark data should not be used outside of research contexts.
@@
We evaluate 6 different LLMs: GPT-4 [3] gpt4-0314, GPT-3.5 [15] gpt-3.5-turbo-0613, both accessed securely through the Azure OpenAI API. We also evaluate LLaMA-2 [7B,70B] [2], Mistral-7B [16], and Mixtral (8x7B) [17]. We evaluate each LLM both for generating summaries and for scoring summaries. We compare methods for scoring with three popular metrics: BERT-Score [18], and ROUGE-1/ROUGE-L [19].
Figure 2: Distribution of scores across different attributes. Attribute Structuring enables auditing which aspects of discharge summaries are generally accurate (e.g. Discharge test) and which are often inaccurate (e.g. Hospital consults).
@@
On a subset of our benchmark, three blinded human annotators evaluate the similarities between the extracted text for each attribute. They score 17 attributes for 30 total documents (20 test summaries generated by GPT-4 and 10 test summaries generated by Mixtral), for a total of 510 attribute labels. Each label is scored on a 4-point scale: Not similar, Somewhat similar, Very similar, Essentially the same. We evaluate the match between the manually annotated attribute labels and the automated attribute labels using Pearson correlation, Spearman (rank) correlation, and root mean-squared-error (RMSE).
@@ 3 Results
ROUGE-1 0.17 / 0.17 0.36 / 0.64 0.34 / 0.48
ROUGE-L 0.26 / 0.18 0.35 / 0.59 0.32 / 0.44
BERT-Score 0.48 / 0.18 0.17 / 0.71 0.23 / 0.69
GPT-4 0.18 / 0.07 0.46 / 0.84 0.38 / 0.77
Table 2: Automatic evaluation scores using different LLMs for generating summaries and for evaluating. The Benchmark column uses Mixtral for structuring summaries as well as evaluating them, making it fully reproducible using open source tools. The remaining columns use GPT-4 for structuring and different LLMs for scoring the attributes. The relative rankings of different LLMs generators is preserved across different columns.
Benchmark Scorer (with GPT-4 structuring)
Mixtral (8x7B) AVG LLaMA-2 (7B) Mistral (7B) LLaMA-2 (70B) Mixtral (8x7B) GPT-3.5 GPT-4
Generator LLaMA-2 (7B) 36.2 45.97 65.4 47.1 49.0 38.5 40.3 35.5
Mistral 37.4 46.37 66.1 47.7 49.2 38.7 40.8 35.7
LLaMA-2 (70B) 38.4 49.70 66.7 50.6 54.3 43.4 43.6 39.6
Mixtral (8x7B) 39.2 51.30 68.4 52.4 54.9 46.0 44.9 41.2
GPT-3.5 39.4 52.78 73.1 53.7 56.7 45.9 45.7 41.6
GPT-4 45.1 56.02 73.8 56.9 61.2 51.2 48.8 44.2
Table 1 shows how well different scoring methods match the annotations provided by humans. Attribute Structuring yields a considerable improvement for all metrics (except for ROUGE-1 when measured with RMSE). Scoring with GPT-4 yields the best match with human annotators. The Pearson correlation coefficient of 0.84 for GPT-4 closely approaches the mean inter-annotator Pearson correlation of 0.90, and the Spearman correlation coefficient of 0.77 similarly nears the mean inter-annotator Spearman correlation of 0.87. The enhanced correlations across all metrics substantiate that attribute-based, fine-grained scoring outperforms holistic methodologies.
Fig. 2 shows how similarities break down across attributes for different attributes / metrics. Some attributes, such as Discharge test are more often correctly identified in the test summaries, whereas others, such as Hospital consults are rarely correct. Aggregating these attribute evaluations through AS allows for improved evaluation compared to evaluating an entire summary at once.
Summaries from stronger LLMs yield higher AS benchmark scores
We now evaluate how well different LLMs perform under an automatic evaluation using Attribute Structuring. The first column of Table 2 shows the benchmark scores achieved by different models; the benchmark scores rely only on the open-source Mixtral model to perform both structuring and scoring, enabling full reproducibility. GPT-4 achieves the highest score, followed by GPT-3.5 and Mixtral (8x7B).
We additionally evaluate how consistent the benchmark scores are with scores provided using different language models for structuring and scoring. The right section of Table 2 shows evaluation scores when structuring with GPT-4 and scoring with various different LLMs. We find that the absolute scores in different columns vary widely, but the relative ordering of LLM generators is almost perfectly preserved between the benchmark and all other settings (the only exception is that the Mixtral scorer slightly prefers the Mixtral generator to the GPT-3.5 generator).
@@
Going beyond aggregate interpretations, Fig. 3 shows an example of how similarity can be audited for a single pair of summaries. Unlike posthoc feature importance [20, 21], or intrinsically interpretable models [22, 23], our interpretable grounding comes directly from an LLM, similar to recent works that use LLMs to generate explanations [24, 25] and ground those explanations in evidence [26, 27]. Since AS decomposes similarity into a series of attributes, the attributes can each be audited by directly prompting an LLM (here, GPT-4) to find a text span that serves as evidence for the extracted attribute in each document. Fig. 3 shows text spans for two discrepancies: Admission diagnosis and Pertinent lab tests, which are only partially retained in the test summary.
@@ 4 Discussion
Attribute Structuring constitutes an important step towards safely deploying LLMs in healthcare settings. As LLMs continue to generally improve, their ability to produce and evaluate clinical summaries with AS is likely to improve as well. Going forward, Attribute Structuring can be harnessed in a variety of ways to improve clinical NLP summarization, e.g. for evaluating free-form text such as conversations [28], improving model distillation by helping to filter effective summaries [29, 30], and improving the explainability of the process [31].
Limitations
One limitation of Attribute Structuring is that it incurs a high computational cost as multiple LLM calls are chained together; however, these costs may continue to decrease as models become more efficient [32]. Another limitation is that LLMs (and thereby Attribute Structuring) continue to be sensitive to prompts, increasing the need for methods to make LLMs more amenable to prompting [15, 33, 34] and to make finding strong prompts easier [35, 36, 37, 38]. Additionally, AS relies on a human-given ontology for evaluation, but could be made more general by developing methods to automatically define extract and define an ontology for a given domain. While Attribute Structuring helps to improve the automatic evaluation of summaries, there are nevertheless risks present with fully relying on LLMs without human supervision.
@@ 1. Generation prompt:
Prompt = "You are tasked with generating a high quality clinical discharge summary for the provided input text INPUT_DOC. The summary has to be very relevant to the input document and cover the most important aspects of the input. Follow the following steps:
@@
Step 5: Return the final discharge summary with all the remaining sections that have contents. Remember to remove sections with no information. Context : "
@@ 2. Structuring prompt:
= ResponseSchema(name = ’main_diag’, description=’diagnosis mostly accountable for the largest portion of the patient"s stay, responsible for the greatest part of the length of stay ’)
@@
prompt = "You are an expert in information extraction and structuring from clinical notes. Given a clinical note, create a structured output. For a given variable, if you can not determine/find a value, return NONE. Dont add any extra text, just the structured value. Here is the note: INPUT_DOC. format_instructions "
@@ 3. Structured attribute Scoring Prompt:
prompt = "You will be given a python dictionary containing a clinical variable as a key and a list containing two values for the variable as values.
@@
Please make sure you read and understand these instructions carefully. Return the similarity score. Return only the score, don’t include any other extra text. "
@@ 4. Document scoring without structuring prompt
prompt = """You will be given two clinical discharge summaries, summary1 and summary2.
Your task is to rate how similar the two summaries are from 1-10. 1 is not similar while 10 is essentially identical. Focus on the following important clinical variables when performing the comparison: How similar is the diagnosis, how similar are the goals, how similar is hospital course and history, how similar are the medications administered, how similar are the physical condition diagnoses, how similar are the followup consults and procedures, how similar are the lab tests performed, how similar are the patients discharge status, is the follow-up instructions similar, are there any similar appointments, and instructions only return the score, dont add any extra text Here are the inputs discharge summaries: summary1: [value1] summary2: [value2]. """
A.1 Prompts given to human annotators

### v2

@@ Abstract
Summarizing clinical text is crucial in health decision-support and clinical research. Large language models (LLMs) have shown the potential to generate accurate clinical text summaries, but still struggle with issues regarding grounding and evaluation, especially in safety-critical domains such as health. Holistically evaluating text summaries is challenging because they may contain unsubstantiated information. Here, we explore a general mitigation framework using Attribute Structuring (AS), which structures the summary evaluation process. It decomposes the evaluation process into a grounded procedure that uses an LLM for relatively simple structuring and scoring tasks, rather than the full task of holistic summary evaluation. Experiments show that AS consistently improves the correspondence between human annotations and automated metrics in clinical text summarization. Additionally, AS yields interpretations in the form of a short text span corresponding to each output, which enables efficient human auditing, paving the way towards trustworthy evaluation of clinical information in resource-constrained scenarios.
keywords:
Large language model, Summarization, Structuring, Medical
Data and Code Availability
We release our full code at https://github.com/microsoft/attribute-structuring. All analyzed data comes from the freely available discharge notes in MIMIC-III (Johnson et al., 2016).
@@ 1 Introduction
Automatic medical text summarization has emerged as a critical task for real-world impact in natural-language processing amidst widespread physician fatigue and burnout. Meanwhile, pre-trained large language models (LLMs) have shown impressive proficiency in a range of complex natural language processing tasks (Brown et al., 2020; dubey2024llama; OpenAI, 2023), including in text summarization (Liu et al., 2022; Zhang et al., 2024). However, evaluating these summaries is challenging, as they rely on human effort to track different attributes and facts.
LLMs may be able to help automate this evaluation. LLMs have previously been used to provide practical evaluations in a variety of evaluation settings (Liu et al., 2023; Fu et al., 2023). However, these evaluations have been imperfect: yielding relatively low correlation with human judgments, showing bias (Zheng et al., 2023; Wang et al., 2023), or more generally struggling with long-form nuanced evaluation, as is often the case in summarization (Zhong et al., 2022).
@@
To address these issues, structuring the scoring process can help to ground the evaluation, and help an LLM decompose a complex evaluation process into simpler, more accurate steps. Following this intuition, we introduce a simple structuring process, which we call Attribute Structuring (AS) (Fig. 1). Given a groundtruth summary and a summary to be evaluated, AS leverages an LLM to score a summary through prompting. Rather than directly asking the LLM for a score, we ask it to provide a score for each pair of attributes, one from the groundtruth and another from the summary to be evaluated. We extract 16 attributes derived from a clinical ontology for each summary; the attributes are shown to define the elements of quality in the context of a clinical discharge summary in a large scale study (Savvopoulos et al., 2018).
Unlike previous methods such as FActScore (Min et al., 2023), AS does not require a reliable knowledge source to verify each attribute, but instead relies on the pre-specified ontology to capture relevant aspects of a summary. DocLens  (Xie et al., 2024), similar to our work, performs a fine-grained evaluation of medical text using various generation models and shows that such scoring has higher agreement with the judgements of medical experts. However, DocLens relies on the model to generate the important attributes rather than a clinical ontology.
We perform experiments on clinical summarization tasks using discharge notes from MIMIC-III (Johnson et al., 2016) and various LLMs. We find that that Attribute Structuring helps reduce the gap between automated metrics and human annotators. In addition, AS enables auditing the evaluation for a particular summary by providing grounding for each attribute, in the form of a short text span in the input.
@@ 2.1 Methods: Attribute Structuring
Fig. 1 shows the various steps of the Attribute Structuring (AS) pipeline. The pipeline takes in two summaries: a groundtruth summary and a generated summary (which we are trying to evaluate). Additionally, it uses a set of attributes that are important for the task at hand. These attributes comes from an ontology that utilizes domain knowledge to define the elements of quality, in our case the context of a clinical discharge summary in a large scale study (Savvopoulos et al., 2018), e.g. Admission diagnosis or Goals of care.
Each step in the pipeline is achieved by prompting an LLM. In the first step, the Attribute Structuring pipeline first extracts text for each attribute in the ontology from both summaries. Then, for each pair of extracted attributes, it generates a similarity score by prompting an LLM to compare them on a scale of 1–4 (all prompts are provided in Appendix A).
@@ 2.2 Experimental setup
Dataset
We analyze discharge summaries for 300 expert selected MIMIC-III notes  (Johnson et al., 2016) and release scripts to process and use this data. The notes are selected based on the quality of their discharge summaries, where quality is defined as having most of the important attributes and contain patient data that can be verified from the clinical notes. To avoid context size bias, we select MIMIC notes with no more than 4,000 tokens (measured with tiktoken’s p50_base tokenizer [^1] https://github.com/openai/tiktoken).
@@
We evaluate 6 different LLMs: GPT-4o (OpenAI, 2023) gpt4-0314, GPT-3.5 (Ouyang et al., 2022) gpt-3.5-turbo-0613, all accessed securely through the Azure OpenAI API. We also evaluate several instruction-finetuned models: LLaMA-2 70B (Touvron et al., 2023), and LLaMA-3.1 [8B, 70B] (dubey2024llama). We compare methods for scoring with two popular metrics: BERT-Score (Zhang et al., 2019), and ROUGE-L (Lin, 2004).
Figure 2: Distribution of scores across different attributes. Attribute Structuring enables auditing which aspects of discharge summaries are generally accurate (e.g. Main diagnosis) and which are often inaccurate (e.g. Lab test).
@@
On a subset of our selected notes, five blinded human annotators (all are expert medical coders) evaluate the similarities between the extracted text for each attribute. They score 16 attributes for 24 total documents (12 test summaries generated by GPT-4 and 12 test summaries generated by LLaMA-3 70B), for a total of 384 attribute labels. Each label is scored on a 4-point scale: Not similar:attributes don’t share similar elements, Somewhat similar: attributes share a few elements, Very similar: attributes share most elements, Essentially the same: attributes share all important elements. We evaluate the match between the manually annotated attribute labels and the automated attribute labels using Pearson correlation, Spearman (rank) correlation, root mean-squared-error (RMSE), and assess inter-annotator agreement with Fleiss’ Kappa (fleiss1971measuring).
@@ 3 Results
Table 1 shows how well different scoring methods match the annotations provided by humans (the inter-annotator Fleiss’ Kappa score is 0.71). Attribute Structuring yields a considerable improvement for all metrics. Scoring with GPT-4 yields the best match with human annotators both in Pearson and Spearman correlations and has the least RMSE error. Even a traditional scoring metric, ROUGE-L, shows significant improvement when attributes are structured before scoring.
The enhanced correlations across all metrics substantiate that attribute-based, fine-grained scoring outperforms holistic methodologies.
Fig. 2 shows the breakdown of similarities across attributes for different metrics. Some attributes, such as Main diagnosis are more often correctly identified in the test summaries, whereas others, such as Lab test and Discharge test are rarely correct. Aggregating these attribute evaluations through AS allows for improved evaluation compared to evaluating an entire summary at once.
@@
ROUGE-L 0.17 / 0.12 0.48 / 0.70 0.36 / 0.70
BERT-Score 0.52 / 0.22 0.13 / 0.70 0.28 / 0.80
GPT-4 1.10 / 0.06 0.54 / 0.91 0.43 / 0.81
Table 2: Automatic evaluation scores using different LLMs for generating summaries.
Summarizer AS Score
LLaMA-2 (70B) 82.8
GPT-3.5 90.0
LLaMA-3.1 (8B) 91.8
GPT-4-0314 93.4
GPT-4o 95.3
LLaMA-3.1 (70B) 95.7
Analysis: AS scores can be used to evaluate the quality of summaries
In settings where there aren’t manual annotations, we can use AS score to evaluate the quality of summaries generated by LLMs. In Table 2, we can see that AS scores for summaries generated by six LLMs are consistent with the expected quality of the individual models.
@@
Going beyond aggregate interpretations, Fig. 3 shows an example of how similarity can be audited for a single pair of summaries. Unlike posthoc feature importance (Lundberg and Lee, 2017; Ribeiro et al., 2016), or intrinsically interpretable models (Rudin et al., 2021; Singh et al., 2023), our interpretable grounding comes directly from an LLM, similar to recent works that use LLMs to generate explanations (MacNeil et al., 2022; Chen et al., 2024) and ground those explanations in evidence (Xue et al., 2023; Gero et al., 2023). Since AS decomposes similarity into a series of attributes, the attributes can each be audited by directly prompting an LLM (here, GPT-4) to find a text span that serves as evidence for the extracted attribute in each document. Fig. 3 shows text spans for two discrepancies: Admission diagnosis and Pertinent lab tests, which are only partially retained in the test summary.
@@ 4 Discussion
Attribute Structuring constitutes an important step towards safely deploying LLMs in healthcare settings. As LLMs continue to generally improve, their ability to produce and evaluate clinical summaries with AS is likely to improve as well. Going forward, Attribute Structuring can be harnessed in a variety of ways to improve clinical NLP summarization, e.g. for evaluating free-form text such as conversations (Tu et al., 2024) or improving model distillation by helping to filter effective summaries (Wu et al., 2023; Toma et al., 2023).
One limitation of Attribute Structuring is that it incurs a high computational cost as multiple LLM calls are chained together; however, these costs may continue to decrease as models become more efficient (Dao et al., 2022). Another limitation is that LLMs (and thereby Attribute Structuring) continue to be sensitive to prompts, increasing the need for methods to make LLMs more amenable to prompting (Ouyang et al., 2022; Zhou et al., 2023; Zhang et al., 2023) and to make finding strong prompts easier (Shin et al., 2020; Singh et al., 2022; Xu et al., 2023; Morris et al., 2023). Additionally, AS relies on a human-given ontology for evaluation, but could be made more general by developing methods to automatically define extract and define an ontology for a given domain.
@@ Appendix A Appendix
Table A1: Automatic evaluation scores using different LLMs for generating and evaluating summaries. All use GPT-4 for structuring and different LLMs for scoring the attributes. The relative rankings of different LLMs generators is preserved across different columns. Results are for a subset of 30 randomly chosen notes.
Scorer (with GPT-4 structuring)
AVG LLaMA-2 (7B) Mistral (7B) LLaMA-2 (70B) Mixtral (8x7B) GPT-3.5 GPT-4
Generator LLaMA-2 (7B) 45.97 65.4 47.1 49.0 38.5 40.3 35.5
Mistral 46.37 66.1 47.7 49.2 38.7 40.8 35.7
LLaMA-2 (70B) 49.70 66.7 50.6 54.3 43.4 43.6 39.6
Mixtral (8x7B) 51.30 68.4 52.4 54.9 46.0 44.9 41.2
GPT-3.5 52.78 73.1 53.7 56.7 45.9 45.7 41.6
GPT-4 56.02 73.8 56.9 61.2 51.2 48.8 44.2
Table A2: Automatic evaluation scores using different LLMs for structuring summaries yield similar rankings between generator models, although GPT-4 structuring generally yields higher scores. Both use Mixtral (8x7B) for scoring. Results are for a subset of 30 randomly chosen notes.
Generator Mixtral structuring GPT-4 structuring
LLaMA-2 (7B) 36.2 38.5
Mistral 37.4 38.7
LLaMA-2 (70B) 38.4 43.4
Mixtral (8x7B) 39.2 46.0
GPT-3.5 39.4 45.9
GPT-4 45.1 51.2
A.1 Prompts
@@ 1. Generation prompt:
Prompt = “You are tasked with generating a high quality clinical discharge summary for the provided input text INPUT_DOC. The summary has to be very relevant to the input document and cover the most important aspects of the input. Follow the following steps:
@@
Step 5: Return the final discharge summary with all the remaining sections that have contents. Remember to remove sections with no information. Context : ”
@@ 2. Structuring prompt:
= ResponseSchema(name = ’main_diag’, description=’diagnosis mostly accountable for the largest portion of the patient”s stay, responsible for the greatest part of the length of stay ’)
@@
prompt = “You are an expert in information extraction and structuring from clinical notes. Given a clinical note, create a structured output. For a given variable, if you can not determine/find a value, return NONE. Dont add any extra text, just the structured value. Here is the note: INPUT_DOC. format_instructions ”
@@ 3. Structured attribute Scoring Prompt:
prompt = “You will be given a python dictionary containing a clinical variable as a key and a list containing two values for the variable as values.
@@
Please make sure you read and understand these instructions carefully. Return the similarity score. Return only the score, don’t include any other extra text.”
@@ 4. Document scoring without structuring prompt
prompt = “You will be given two clinical discharge summaries, summary1 and summary2.
Your task is to rate how similar the two summaries are from 1-10. 1 is not similar while 10 is essentially identical. Focus on the following important clinical variables when performing the comparison: How similar is the diagnosis, how similar are the goals, how similar is hospital course and history, how similar are the medications administered, how similar are the physical condition diagnoses, how similar are the followup consults and procedures, how similar are the lab tests performed, how similar are the patients discharge status, is the follow-up instructions similar, are there any similar appointments, and instructions only return the score, dont add any extra text Here are the inputs discharge summaries: summary1: [value1] summary2: [value2].”
A.2 Prompts given to human annotators